The price of a stock, which pays no dividends, is $30 and the strike price of a one year European call option on the stock is $25. The risk-free rate is 4% and is continuously compounded. Which of the following is a lower bound for the option such that there are arbitrage opportunities if the price is below the lower bound and no arbitrage opportunities if it is above the lower bound?

$5.98

$3.98

$6.98

$4.98
Answer #1

Time years =1
Stock Price =30
Strike Price =25
Lower Bound =Stock Price -Strike Price*e^(-rt) =30-25*e^(-4%*1) =5.98


Option a is correct option